# Synesthesia of Machines (SoM)-Enhanced ISAC Precoding for Vehicular Networks with Double Dynamics

### v1

@@ III-B2 Optimizing $\bm{F}_{\text{BB,m}}$
$\underset{p_{\text{u,m}}}{\text{min}}$ $~{}\sum_{m=1}^{M}\sum_{u=1}^{U}\|\bm{f}_{\text{m,u}}-p_{\text{m,u}}\bm{f}_{\text{RF,u}}\|_{2}^{2}$ (20)
s.t. $~{}\sum_{m=1}^{M}\sum_{u=1}^{U}p_{\text{m,u}}^{2}=\frac{1}{N_{\text{t}}},$ (20a)
where $\bm{f}_{\text{RF,u}}=\bm{F}_{\text{RF}}[:,u]$. By briefly dropping the power constraint, $p_{\text{m,u}}$ has the close-form solution through least square (LS) method as
$\hat{p}_{\text{m,u}}=\bm{f}_{\text{RF,u}}^{\dagger}\bm{f}_{\text{m,u}},\vspace{-0.2cm}$ (21)
where $\bm{f}_{\text{RF,u}}^{\dagger}=\frac{1}{N_{\text{t}}}\bm{f}_{\text{RF,u}}^{\text{H}}$ is the pseudo inverse of $\bm{f}_{\text{RF},u}$. Taking into account power budget, $\hat{p}_{\text{m,u}}$ can be further normalized as
$p_{\text{m,u}}=\frac{1}{\sqrt{\sum_{m=1}^{M}\sum_{u=1}^{U}\hat{p}_{\text{m,u}}^{2}}}\hat{p}_{\text{m,u}}.\vspace{-0.2cm}$ (22)
@@ III-C Remark on optimization-based schemes
  • Mediocre performance: The obtained solutions may incur potentially significant suboptimality for two main reasons. Firstly, the relaxation of the rank-1 constraint can lead to being stuck in local optimality when optimizing $\bm{F}_{\text{m}}$. Secondly, the decomposition of $\bm{F}_{\text{m}}$ introduces additional loss in overall performance.
@@ IV-2 Stage II: Data transmission
Various frame structures that might be utilized in doubly-dynamic ISAC scenarios are compared, as depicted in Fig. 2. The operation in [34] conducts the precoding design based on the real-time channel estimation and target detection through downlink pilots and uplink feedback in each subframe. This leads to a decrease in SE due to high time overhead and is not suitable for double dynamics. In contrast, the frame structure in [35, 19] aims to predict CSI and the target’s position proactively to reduce prior acquisition overhead. However, real-time optimization of precoding remains essential. The proposed ISAC frame structure mitigates the reliance on the real-time prior, greatly reducing the signal overhead. Moreover, precoding is updated without optimization from scratch, lowering the computational complexity.
@@ V-A Environmental State Space
where $\mathcal{H}^{(n)}$, $\mathcal{U}^{(n)}$ and $\mathcal{T}^{(n)}$ denote the instantaneous state of channel information, users’ positions and the target’s position, respectively. $\mathcal{H}^{(n)}=[\bm{H}_{\text{0}}^{(n)},\cdots,\bm{H}_{N_{d}-1}^{(n)}]$ contains the time-domain channel in $N_{\text{d}}$ delay taps. $\mathcal{U}^{(n)}=\{(\theta_{\text{u}}^{(n)},d_{\text{u}}^{(n)})|_{u=1}^{U}\}$ and $\mathcal{T}^{(n)}=\{(\theta_{\text{s}}^{(n)},d_{\text{s}}^{(n)})\}$ contain the polar coordinates of the users and of the target in the $n$-th subframe, respectively.
@@ V-A2 Position-related Observation
where $(\hat{\theta}_{u}^{(n)},\hat{d}_{u}^{(n)})$ and $(\hat{\theta}_{s}^{(n)},\hat{d}_{s}^{(n)})$ are the estimated coordinates of the users and the target respectively. Denote $\mathcal{P}_{\text{ob}}^{(n)}=\mathcal{U}_{\text{ob}}^{(n)}\cup\mathcal{T}_{\text{ob}}^{(n)}=\{({\theta_{i}^{\text{ob}}}^{(n)},{d_{i}^{\text{ob}}}^{(n)})|_{i=1}^{U+1}\}$ as the set of the estimated coordinates of the objects through echo processing in the $n$-th subframe. Note that BS itself lacks the ability to differentiate between $\mathcal{U}_{\text{ob}}^{(n)}$ and $\mathcal{T}_{\text{ob}}^{(n)}$ from $\mathcal{P}_{\text{ob}}^{(n)}$. In fact, users can periodically upload their locations from GPS to the BS every $T_{\text{fb}}$ subframes, as operated in [36]. These feedbacks are denoted by $\{({\theta_{u}^{\text{fb}}}^{(n)},{d_{u}^{\text{fb}}}^{(n)})_{u=1}^{U}\}$. In subframes with the real-time uploaded feedback, the BS differentiates $U$ users from $\mathcal{P}_{\text{ob}}^{(n)}$ in sequence, based on the nearest neighbor criterion:
@@
$\bm{P}^{(n)}(n_{\text{x}},\!n_{\text{y}})\!=\!\!\begin{cases}1,\!\!&\text{if $\{[\theta_{n_{\text{x}}},\!\theta_{n_{\text{x}}\!+\!1})\!\times\![d_{n_{\text{y}}},\!d_{n_{\text{y}}\!+\!1})\}\!\cap\mathcal{U}_{\text{ob}}^{(n)}\!\!\neq\!\varnothing$},\\ 2,\!\!&\text{if $\{[\theta_{n_{\text{x}}},\!\theta_{n_{\text{x}}\!+\!1})\!\times\![d_{n_{\text{y}}},\!d_{n_{\text{y}}\!+\!1})\}\!\cap\mathcal{T}_{\text{ob}}^{(n)}\!\!\neq\!\varnothing$},\\ 0,\!\!&\text{Otherwise},\end{cases}\vspace{-0.2cm}$ (29)
@@ V-B1 Action Space Design
The hybrid action combination can be expressed as $\bm{A}^{(n)}=\{\bm{a}^{(n)}_{\text{RF},1},\bm{a}^{(n)}_{\text{RF},2},\bm{a}^{(n)}_{\text{BB}}\}$. $\bm{a}_{\text{RF},1}^{(n)}\in[0,1]^{\text{U}^{\text{max}}}$ represents for the probability distribution of the $U$ columns in $\bm{F}_{\text{RF}}^{(n-1)}$ to be selected, and $\bm{a}_{\text{RF},2}^{(n)}\in[0,1]^{N_{\text{t}}}$ represents for the incremental update of the $N_{\text{t}}$ elements in the selected column. $\bm{a}^{(n)}_{\text{BB}}\in[-1,1]^{\text{M}}$ denotes the increment of $p_{\text{m,u}}$ in $\bm{F}_{\text{BB},m}$ ($m\in[1,M]$).
@@
In this updating scheme, the action is defined as $\bm{A}^{(n)}=[\bm{a}_{\text{RF}}^{(n)},\bm{a}_{\text{BB}}^{(n)}]$. $\bm{a}_{\text{RF}}^{(n)}\in[0,1]^{N_{\text{t}}}$ denotes the probability distribution of the $N_{\text{t}}$ antennas being selected to be updated in the $n$-th subframe. $\bm{a}_{\text{BB}}^{(n)}\in[-1,1]^{\text{M}}$ denotes the phases of the differentials at $M$ subcarriers for digital update.
@@ V-D1 Transition of $\mathcal{U}^{(n)}$ and $\mathcal{T}^{(n)}$
$\sin{(\theta_{\text{u}}^{(n+1)}\!-\!\theta^{(n)})}=|\bm{v}_{\text{u}}^{(n)}|\Delta T\sin{(\theta_{\text{u}}^{(n)}\!-\!\angle(\bm{v}_{\text{u}}^{(n)})\!+\!\pi)}/d_{\text{u}}^{(n+1)},$
$(d_{\text{u}}^{(n+1)})^{2}\!-\!(d_{\text{u}}^{(n)})^{2}=(|\bm{v}_{\text{u}}^{(n)}|\Delta T)^{2}$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}-2d_{\text{u}}^{(n)}|\bm{v}_{\text{u}}^{(n)}|\Delta T\cos{(\theta_{\text{u}}^{(n)}-\angle(\bm{v}_{\text{u}}^{(n)})\!+\!\pi)}.\vspace{-0.2cm}$ (41)
The target position is $\mathcal{T}^{(n+1)}=\{(\theta_{\text{s}}^{(n+1)},d_{\text{s}}^{(n+1)})\}$, evolving similarly according to $\bm{v}_{\text{s}}^{(n)}$.
@@ V-D2 Transition of $\mathcal{H}^{(n)}$
where $\Delta\beta\sim\mathcal{CN}(0,\sigma_{\Delta\beta})^{2}$ denote the variation of the complex gain of the paths. At BS, the channel-related observation $\bm{S}_{\text{H}}^{(n+1)}=\bm{S}_{\text{H}}^{(n+1)}$ remains the initial estimate, while the position-related observation $\bm{S}_{\text{P}}^{(n+1)}=\left\{\bm{P}^{(n-2)},\bm{P}^{(n-1)},\bm{P}^{(n)}\right\}$ is updated with the latest estimation, $\bm{P}^{(n)}$, through echo processing in the $n$-th subframe.
Figure 4: An illustration of the environmental state transition.
@@ V-E5 Value decoder
In this PSAC, the actor network $\mu$, parameterized by $\phi_{\text{A}}=\phi_{\text{SE}}\cup\phi_{\text{I}}\cup\phi_{\text{AD}}$, maps from the state space to an action $\bm{A}\in\mathcal{A}$ as $\bm{A}^{(n)}=\mu(\mathcal{O}^{(n)}|\phi_{\text{A}})$. The critic network, parameterized by $\phi_{\text{C}}=\phi_{\text{SE}}\cup\phi_{\text{AE}}\cup\phi_{\text{IN}}\cup\phi_{\text{VD}}$, maps from the joint observation space and action space into reward, playing a role of action evaluation as $Q(\mathcal{O}^{(n)},\bm{A}^{(n)}|\phi_{\text{C}})$. The PSAC network acts as actor when the green area is activated, and critic when the pink area is activated, with the brown area being the shared part. The overall parameters are expressed as $\phi=\phi_{\text{A}}\cup\phi_{\text{C}}$. In the offline training stage, the transition tuple $(\mathcal{O}^{(n)},\bm{A}^{(n)},r^{(n)},\mathcal{O}^{(n+1)})$ is stored in a memory replay buffer $\mathcal{M}$ for network updating.
@@
1:Initialize PSAC network with random weights $\phi\!=\!\phi_{\text{A}}\!\cup\phi_{\text{C}}$;
@@
8:         Select action $\bm{a}^{(n)}=\mu(\bm{o}^{(n)}|\phi_{\text{A}})$, update $\bm{F}_{\text{RF}}$ and $\bm{F}_{\text{BB},m}$ based on $\bm{a}^{(n)}$;
@@
14:         Sample a batch of $N_{b}$ tuples from $\mathcal{M}$ randomly;
@@
17:         Update critic loss as $L_{\text{c}}^{(k)}\!=\!L_{\text{c}}^{(k-1)}\!+\!\frac{1}{N_{b}}\sum_{i=1}^{N_{b}}\delta_{i}$;
@@ V-F Network Update
The use of shared parameters in the PSAC architecture introduces an additional challenge in training stability compared to traditional actor-critic architectures. To address this challenge and enhance training stability, we have drawn inspiration from a previous study [38] and implemented an adaptive loss function specifically designed for the PSAC network. To reduce fluctuations in target values during training and expedite convergence, we employ target networks $Q^{\prime}$ and $\mu^{\prime}$ parameterized by $\phi_{\text{C}}^{\prime}$ and $\phi_{A}^{\prime}$ respectively. During each episode, $K$ batches are sampled successively from $\mathcal{M}$ for network update. When the $k$-th batch $(k=1,\cdots,K)$ of $N_{b}$ transition tuples, $\{(\bm{o}_{i}^{(k)},\bm{a}_{i}^{(k)},r_{i}^{(k)},\bm{o}_{i+1}^{(k)})|_{i=1}^{N_{\text{b}}}\}$, are randomly sampled, the target reward of the $i$-th sample ($i=1,\cdots,N_{b}$) is calculated as
@@
TABLE II: System setup for environment generation

### v2

@@ III-B2 Optimizing $\bm{F}_{\text{BB},m}$
$\underset{p_{u,m}}{\text{min}}$ $~{}\sum_{m=1}^{M}\sum_{u=1}^{U}\|\bm{f}_{m,u}-p_{m,u}\bm{f}_{\text{RF},u}\|_{2}^{2}$ (20)
s.t. $~{}\sum_{m=1}^{M}\sum_{u=1}^{U}p_{m,u}^{2}=\frac{1}{N_{\text{t}}},$ (20a)
where $\bm{f}_{\text{RF},u}=\bm{F}_{\text{RF}}[:,u]$. By briefly dropping the power constraint, $p_{m,u}$ has the close-form solution through least square (LS) method as
$\hat{p}_{m,u}=\bm{f}_{\text{RF},u}^{\dagger}\bm{f}_{m,u},\vspace{-0.2cm}$ (21)
where $\bm{f}_{\text{RF},u}^{\dagger}=\frac{1}{N_{\text{t}}}\bm{f}_{\text{RF},u}^{\text{H}}$ is the pseudo inverse of $\bm{f}_{\text{RF},u}$. Taking into account power budget, $\hat{p}_{m,u}$ can be further normalized as
$p_{m,u}=\frac{1}{\sqrt{\sum_{m=1}^{M}\sum_{u=1}^{U}\hat{p}_{m,u}^{2}}}\hat{p}_{m,u}.\vspace{-0.2cm}$ (22)
@@ III-C Remark on optimization-based schemes
  • Mediocre performance: The obtained solutions may incur potentially significant suboptimality for two main reasons. Firstly, the relaxation of the rank-1 constraint can lead to being stuck in local optimality when optimizing $\bm{F}_{m}$. Secondly, the decomposition of $\bm{F}_{m}$ introduces additional loss in overall performance.
@@ IV-2 Stage II: Data transmission
Various frame structures that can be utilized in doubly-dynamic ISAC scenarios are compared, as depicted in Fig. 2. The operation in [38] conducts the precoding design based on the real-time channel estimation and target detection through downlink pilots and uplink feedback in each subframe. This leads to a decrease in SE due to high time overhead and computational burden, making it unsuitable for double dynamics. In contrast, the frame structure in [39, 20] proactively predicts CSI and the target’s position to minimize the time overhead associated with prior acquisition. However, real-time optimization of precoding remains crucial, and the computational complexity remains essential. The proposed ISAC frame structure alleviates the dependence on real-time prior acquisition, significantly reducing time overhead. Moreover, precoding updates are performed without the need for complete optimization, thus lowering computational complexity.
@@ V-A Environmental State Space
where $\mathcal{H}^{(n)}$, $\mathcal{U}^{(n)}$ and $\mathcal{T}^{(n)}$ denote the instantaneous state of channel information, users’ positions and the target’s position, respectively. $\mathcal{H}^{(n)}=[\bm{H}_{0}^{(n)},\cdots,\bm{H}_{N_{d}-1}^{(n)}]$ contains the time-domain channel in $N_{\text{d}}$ delay taps. $\mathcal{U}^{(n)}=\{(\theta_{u}^{(n)},d_{u}^{(n)})|_{u=1}^{U}\}$ and $\mathcal{T}^{(n)}=\{(\theta_{\text{s}}^{(n)},d_{\text{s}}^{(n)})\}$ contain the polar coordinates of the users and of the target in the $n$-th subframe, respectively.
@@ V-A1 Channel-related Observation
TABLE II: Notations related to the environmental observations.
Notation Definition
$\bm{S}_{\text{H}}^{(n)}$ Channel-related observations as network input
$\bm{S}_{\text{P}}^{(n)}$ Position-related observations as network input
$\widehat{\bm{H}}_{d}^{(0)}$ Estimate of the $d$-th tap of initial CSI
$\bm{D}_{\text{t}}$ Beamspace dictionary
$\mathcal{P}_{\text{ob}}^{(n)}$ The set of coordinates of user-target estimates
$\mathcal{U}_{\text{ob}}^{(n)}$ The BS-end estimates of the users’ positions
$\mathcal{T}_{\text{ob}}^{(n)}$ The BS-end estimates of the target’s position
$({\theta_{i}^{\text{ob}}}^{(n)},{d_{i}^{\text{ob}}}^{(n)})$ BS-end estimation of the $i$-th object’s coordinates
$({\theta_{u}^{\text{fb}}}^{(n)},{d_{u}^{\text{fb}}}^{(n)})$ Periodic GPS feedback from the $u$-th user
@@ V-A2 Position-related Observation
where $(\hat{\theta}_{u}^{(n)},\hat{d}_{u}^{(n)})$ and $(\hat{\theta}_{s}^{(n)},\hat{d}_{s}^{(n)})$ are the estimated coordinates of the users and the target respectively. Denote $\mathcal{P}_{\text{ob}}^{(n)}=\mathcal{U}_{\text{ob}}^{(n)}\bigcup\mathcal{T}_{\text{ob}}^{(n)}=\{({\theta_{i}^{\text{ob}}}^{(n)},{d_{i}^{\text{ob}}}^{(n)})|_{i=1}^{U+1}\}$ as the set of the estimated coordinates of the objects through echo processing in the $n$-th subframe. Note that BS itself lacks the ability to differentiate between $\mathcal{U}_{\text{ob}}^{(n)}$ and $\mathcal{T}_{\text{ob}}^{(n)}$ from $\mathcal{P}_{\text{ob}}^{(n)}$. In fact, users can periodically upload their locations from GPS to the BS every $T_{\text{fb}}$ subframes, as operated in [40]. These feedbacks are denoted by $\{({\theta_{u}^{\text{fb}}}^{(n)},{d_{u}^{\text{fb}}}^{(n)})_{u=1}^{U}\}$. In subframes with the real-time uploaded feedback, the BS differentiates $U$ users from $\mathcal{P}_{\text{ob}}^{(n)}$ in sequence, based on the nearest neighbor criterion:
@@
$\bm{P}^{(n)}(n_{\text{x}},\!n_{\text{y}})\!=\!\!\begin{cases}1,\!\!&\text{if $\{[\theta_{n_{\text{x}}},\!\theta_{n_{\text{x}}\!+\!1})\!\times\![d_{n_{\text{y}}},\!d_{n_{\text{y}}\!+\!1})\}\!\bigcap\mathcal{U}_{\text{ob}}^{(n)}\!\!\neq\!\emptyset$},\\ 2,\!\!&\text{if $\{[\theta_{n_{\text{x}}},\!\theta_{n_{\text{x}}\!+\!1})\!\times\![d_{n_{\text{y}}},\!d_{n_{\text{y}}\!+\!1})\}\!\bigcap\mathcal{T}_{\text{ob}}^{(n)}\!\!\neq\!\emptyset$},\\ 0,\!\!&\text{Otherwise},\end{cases}\vspace{-0.2cm}$ (29)
@@ V-B1 Action Space Design
The hybrid action combination can be expressed as $\bm{A}^{(n)}=\{\bm{a}^{(n)}_{\text{RF},1},\bm{a}^{(n)}_{\text{RF},2},\bm{a}^{(n)}_{\text{BB}}\}$. $\bm{a}_{\text{RF},1}^{(n)}\in[0,1]^{\text{U}^{\text{max}}}$ represents for the probability distribution of the $U$ columns in $\bm{F}_{\text{RF}}^{(n-1)}$ to be selected, and $\bm{a}_{\text{RF},2}^{(n)}\in[0,1]^{N_{\text{t}}}$ represents for the incremental update of the $N_{\text{t}}$ elements in the selected column. $\bm{a}^{(n)}_{\text{BB}}\in[-1,1]^{M}$ denotes the increment of $p_{m,u}$ in $\bm{F}_{\text{BB},m}$ ($m\in[1,M]$).
@@
In this updating scheme, the action is defined as $\bm{A}^{(n)}=[\bm{a}_{\text{RF}}^{(n)},\bm{a}_{\text{BB}}^{(n)}]$. $\bm{a}_{\text{RF}}^{(n)}\in[0,1]^{N_{\text{t}}}$ denotes the probability distribution of the $N_{\text{t}}$ antennas being selected to be updated in the $n$-th subframe. $\bm{a}_{\text{BB}}^{(n)}\in[-1,1]^{M}$ denotes the phases of the differentials at $M$ subcarriers for digital update.
@@ V-B2 Precoding Update
Remark 1.
The updates across the user dimension are compatible with the multi-beam structure (MBS) using lens arrays, while updates across the antenna dimension are suitable for multi-aperture structures (MAS) with PS networks. Therefore, the proposed updating schemes can adapt to existing hybrid precoding hardware architectures.
@@ V-D1 Transition of $\mathcal{U}^{(n)}$ and $\mathcal{T}^{(n)}$
$\sin{(\theta_{u}^{(n+1)}\!-\!\theta_{u}^{(n)})}=|\bm{v}_{u}^{(n)}|\Delta T\sin{(\theta_{u}^{(n)}\!-\!\angle(\bm{v}_{u}^{(n)})\!+\!\pi)}/d_{u}^{(n+1)},$
$(d_{u}^{(n+1)})^{2}\!-\!(d_{u}^{(n)})^{2}=(|\bm{v}_{u}^{(n)}|\Delta T)^{2}$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}-2d_{u}^{(n)}|\bm{v}_{u}^{(n)}|\Delta T\cos{(\theta_{u}^{(n)}-\angle(\bm{v}_{u}^{(n)})\!+\!\pi)}.\vspace{-0.2cm}$ (41)
The target position is $\mathcal{T}^{(n+1)}=\{(\theta_{\text{s}}^{(n+1)},d_{\text{s}}^{(n+1)})\}$, evolving similarly according to $\bm{v}_{\text{s}}^{(n)}$. It can be observed that the positions at $(t+1)$ is determined by the positions and the velocities at $t$. Thus the evolutions of $\mathcal{U}^{(n)}$ and $\mathcal{T}^{(n)}$ are Markov chains.
Figure 4: An illustration of the environmental state transition.
@@ V-D2 Transition of $\mathcal{H}^{(n)}$
where $\Delta\beta\sim\mathcal{CN}(0,\sigma_{\Delta\beta})^{2}$ denote the variation of the complex gain of the paths. The parameters of channels satisfy the Markov property, meaning that $\mathcal{H}^{(n)}$ also forms a Markov chain.
At BS, the channel-related observation $\bm{S}_{\text{H}}^{(n+1)}=\bm{S}_{\text{H}}^{(n+1)}$ remains the initial estimate, while the position-related observation $\bm{S}_{\text{P}}^{(n+1)}=\left\{\bm{P}^{(n-2)},\bm{P}^{(n-1)},\bm{P}^{(n)}\right\}$ is updated with the latest estimation, $\bm{P}^{(n)}$, through echo processing in the $n$-th subframe. It can be verified that the observations also exhibit Markov properties.
@@ V-E5 Value decoder
In this PSAC, the actor network $\mu$, parameterized by $\phi_{\text{A}}=\phi_{\text{SE}}\bigcup\phi_{\text{I}}\bigcup\phi_{\text{AD}}$, maps from the state space to an action $\bm{A}\in\mathcal{A}$ as $\bm{A}^{(n)}=\mu(\mathcal{O}^{(n)}|\phi_{\text{A}})$. The critic network, parameterized by $\phi_{\text{C}}=\phi_{\text{SE}}\bigcup\phi_{\text{AE}}\bigcup\phi_{\text{IN}}\bigcup\phi_{\text{VD}}$, maps from the joint observation space and action space into reward, playing a role of action evaluation as $Q(\mathcal{O}^{(n)},\bm{A}^{(n)}|\phi_{\text{C}})$. The PSAC network acts as actor when the green area is activated, and critic when the pink area is activated, with the brown area being the shared part. The overall parameters are expressed as $\phi=\phi_{\text{A}}\bigcup\phi_{\text{C}}$. Note that the updating schemes across user and antenna dimensions differ slightly in the action encoder and action decoder due to customized hybrid action spaces. For brevity, we illustrate the update across the user dimension in Fig. 5, while the update across the antenna dimension follows a similar pattern.
@@
1:Initialize PSAC network with random weights $\phi\!=\!\phi_{\text{A}}\!\bigcup\phi_{\text{C}}$;
@@
8:         Select action $\bm{a}^{(n)}=\mu(\bm{o}^{(n)}|\phi_{\text{A}})$ with $\xi$-greedy policy, update $\bm{F}_{\text{RF}}$ and $\bm{F}_{\text{BB},m}$ based on $\bm{a}^{(n)}$;
@@
14:         Sample a batch of $N_{\text{b}}$ tuples from $\mathcal{M}$ randomly;
@@
17:         Update critic loss as $L_{\text{c}}^{(k)}\!=\!L_{\text{c}}^{(k-1)}\!+\!\frac{1}{N_{\text{b}}}\sum_{i=1}^{N_{\text{b}}}\delta_{i}$;
@@ V-F Network Update
The use of shared parameters in the PSAC architecture introduces an additional challenge in training stability compared to traditional actor-critic architectures. To address this challenge and enhance training stability, we have drawn inspiration from a previous study [42] and implemented an adaptive loss function specifically designed for the PSAC network. To reduce fluctuations in target values during training and expedite convergence, we employ target networks $Q^{\prime}$ and $\mu^{\prime}$ parameterized by $\phi_{\text{C}}^{\prime}$ and $\phi_{A}^{\prime}$ respectively. In the offline training stage, the transition tuple $(\mathcal{O}^{(n)},\bm{A}^{(n)},r^{(n)},\mathcal{O}^{(n+1)})$ is stored in a memory replay buffer $\mathcal{M}$ for network updating. During each episode, $K$ batches are sampled successively from $\mathcal{M}$ for network update. When the $k$-th batch $(k=1,\cdots,K)$ of $N_{b}$ transition tuples, $\{(\bm{o}_{i}^{(k)},\bm{a}_{i}^{(k)},r_{i}^{(k)},\bm{o}_{i+1}^{(k)})|_{i=1}^{N_{\text{b}}}\}$, are randomly sampled, the target reward of the $i$-th sample ($i=1,\cdots,N_{b}$) is calculated as
@@
Then we set the unified loss function for PSAC as $Z^{(k)}(\phi)=\frac{1}{N_{\text{b}}}\sum_{i=1}^{N_{b}}Z_{i}^{(k)}$, where
$\phi^{(k)}$ $=\phi^{(k-1)}-\kappa\nabla_{\phi}Z^{(k)}(\phi)$ (48)
V-G Convergence Analysis
We then delve into the theoretical analysis of the convergence of the proposed scheme, to guarantee its stability. Note that PSAC is a specialized variant of the actor-critic architecture, with the actor sharing part of parameters with the critic. In fact, the shared parameters and objective functions do not affect the update of their respective networks. Therefore, the convergence analysis of the proposed PSAC is similar to that of the standard actor-critic [43, 44].
Proposition 2.
Denoting $J^{(t)}(\mu)$ as the expected cumulative reward achieved by the actor at the $t$-th iteration. The output of the PSAC in Algorithm 2 satisfies
$\min_{t\in[1,T]}\mathbb{E}\|\nabla J^{(t)}(\mu)\|^{2}\leqslant\frac{c_{1}}{T}+\frac{c_{2}}{N_{\text{b}}}+c_{3}\kappa^{2},\vspace{-0.2cm}$ (49)
where $c_{1}$, $c_{2}$, $c_{3}$ are constants and $\kappa$ denotes the approximation errors of the Q-value from the critic.
Proof.
See Appendix B. ∎
Therefore, the policy gradient approaches zero as the number of episodes becomes sufficiently large. The convergence upper bound of PSAC consists of the convergence rate term vanishing with iterations, the variance term caused by stochastic sampling and controllable by $N_{\text{b}}$, and the system error term composed of approximation errors. In practice, the high capacity of the neural networks can help to reduce the error and achieves better convergence accuracy.
TABLE III: System setup for environment generation